# The Brain’s Bitter Lesson: Scaling Speech Decoding With Self-Supervised Learning

### v3

@@ 4 Experiments
In this section, we evaluate the representations learned with our pretext tasks by measuring their ability to scale downstream performance with unlabelled data. This includes understanding how well they can generalise across datasets, subjects, and tasks. We focus our evaluation on MEG data as the signal is rich, with better spatial resolution than EEG (Lopes da Silva, 2013) and faster sampling rates than fMRI (Hall et al., 2014).
We pre-train all models to completion and then fine-tune on labelled data for each task. In all tables and figures, we quote the receiver operating characteristic area under the curve (ROC AUC) where chance is always 0.5 regardless of class balance. We show the test ROC AUC at the best validation ROC AUC (early stopping) and quote uncertainty as the standard error of the mean over three seeds. Additionally, we state the $t$-score and $p$-value from single-sample one-sided $t$-tests against chance.
@@ 4.1 Experimental setup
Datasets. Unless specified otherwise, our experiments use Cam-CAN (Shafto et al., 2014; Taylor et al., 2017) as an unlabelled representation learning dataset for pre-training. This is a study containing 641 subjects with resting and sensorimotor tasks, totalling approximately 160 hours of MEG recordings. For our downstream tasks, we use two labelled heard speech MEG datasets where participants listen to short stories or audiobooks. Armeni et al. (2022) contains 3 subjects who listen to 10 hours of recordings each (30 hours total) while Gwilliams et al. (2023) has 27 subjects, each recorded for 2 hours (54 hours total). Overall, we utilise over 200 hours of data. To the best of our knowledge, this is the largest volume of MEG data ever used for speech decoding.
Preprocessing. Each recording is in $\mathbb{R}^{S\times T}$ where $S$ is the number of sensors and $T$ is the number of time points sampled by the scanner. To eliminate high-frequency muscle movement artifacts, we apply a low-pass filter at 125Hz as well as a high-pass filter at 0.5Hz to remove slow-drift artifacts. Since the datasets were recorded in Europe, where the electric grid frequency is 50Hz, we apply a notch filter at multiples of 50Hz to account for line noise. Treating the low-pass filter threshold as the Nyquist frequency, we downsample the signal to twice that at 250Hz, avoiding aliasing within our band of interest. Finally, we detect bad sensor channels, those with significant noise and artifacts, using a variance threshold and replace them by interpolating the spatially nearest sensors.
Downstream tasks. We evaluate our methods with two fundamental speech decoding tasks of increasing difficulty. The first, speech detection, determines whether speech occurs in the auditory stimulus using the neural response. The second task is voicing classification. Given data aligned at the occurrence of a phoneme, the task is to recognise whether the phoneme is voiced or voiceless, where voicing is a binary phonetic feature that categorises whether a speech sound is associated with vocal cord vibration. We select these tasks as they are simpler than phoneme recognition, but are foundational because they must be solved to decode speech accurately into natural language.
@@ 4.2 Learning Generalisable Representations Using Pretext Tasks
Our first experiment investigates whether our self-supervised objectives produce generalisable representations. In Table 1, we show the results of pre-training models with each pretext task independently as well as together. Here, all of our pretext tasks lead to results that are statistically significant, and outperform a baseline fine-tuned without pre-training. This provides initial evidence that our tasks are helpful in speech decoding. Interestingly, the combination of all pretext tasks leads to better generalisation than any task on its own. As we hypothesised earlier, this may be because our pretext tasks capture complementary properties in time- and frequency-space, enforcing that our representation includes more salient features for speech decoding than any individual task.
Armeni Gwilliams
Experiment ROC AUC $t$ $p$ ROC AUC $t$ $p$
Linear $0.559\pm 2\mathrm{e}{-4}$ $341$ $4\mathrm{e}{-6}$ $0.527\pm 7\mathrm{e}{-5}$ $379$ $3\mathrm{e}{-6}$
BIOT + linear $0.500\pm 4\mathrm{e}{-4}$ 00 $6\mathrm{e}{-1}$ $0.499\pm 2\mathrm{e}{-4}$ $-3$ $1\mathrm{e}{+0}$
Ours No pre-training $0.519\pm 0.002$ $8$ $7\mathrm{e}{-3}$ $0.498\pm 0.003$ 00 $7\mathrm{e}{-1}$
+ linear Amp($\rho=0.2$) $0.602\pm 0.001$ $114$ $4\mathrm{e}{-5}$ $0.532\pm 0.005$ $6$ $1\mathrm{e}{-2}$
Phase($\rho=0.5$) $0.603\pm 0.003$ $35$ $4\mathrm{e}{-4}$ $0.535\pm 0.003$ $12$ $3\mathrm{e}{-3}$
Band $0.616\pm 0.003$ $44$ $3\mathrm{e}{-4}$ $0.542\pm 0.001$ $46$ $2\mathrm{e}{-4}$
All tasks $\mathbf{0.621}\pm 0.003$ $36$ $4\mathrm{e}{-4}$ $\mathbf{0.543}\pm 0.003$ $13$ $3\mathrm{e}{-3}$
@@
Table 1: Pre-training with pretext tasks leads to better representations for decoding speech. In the linear-only case, we train a supervised linear classifier on the input MEG signals. For BIOT, we train a linear layer on top of the pre-trained BIOT representation, with the rest of the model frozen. Similarly, for ours, we train a linear probe on top of our pre-trained backbone with its weights frozen. In the no pre-training baseline, the backbone uses randomly initialised and subsequently unmodified weights. When all pretext tasks are used, their losses are weighted equally.
Now, we turn to the other baselines. Our approach significantly outperforms the equivalent with a raw MEG input instead of a pre-trained representation (the linear experiment). Here, the baseline has substantially more trainable parameters because the input dimension is far larger without an encoder. Even with this bias favouring the experiment with the raw input, using our representation still performs better. We also compare our approach to BIOT (Yang et al., 2023) which is a similar state-of-the-art self-supervised method. When BIOT is pre-trained using exactly the same data, the fine-tuned probe fails to generalise entirely after exhaustive hyperparameter tuning. We put this down to three critical reasons. Firstly, BIOT was designed around considerably lower-dimensional signals. Their EEG evaluation used an order of magnitude fewer sensors than our MEG data. With MEG, their transformer approach requires many more channel embeddings, leading to difficulty learning the complex interactions between sensors. Secondly, our self-supervised objective extracts speech decoding features which is essential for solving speech decoding tasks. BIOT performs well on simple EEG tasks in Yang et al. (2023)’s evaluation, but non-invasive speech decoding is significantly more challenging. Together, these obstacles suggest a vast amount of data is required to learn their objective with MEG. Indeed, given that they pre-train with over 50 thousand hours of EEG data in their evaluation, their objective appears too general to efficiently learn a representation for speech decoding from the limited amount of MEG pre-training data (160 hours) available to us. This highlights the importance of data-efficiency in SSL methods for MEG.
Among the individual pretext tasks, band prediction leads the rest. Perhaps this is because, by learning to discriminate between meaningful bands, the representation easily identifies phase-locking to speech onset in theta waves (Peelle et al., 2012). Further investigation is necessary here. The choice of the proportion of sensors to apply transformations to, $\rho=0.5$ for phase shift prediction and $\rho=0.2$ for amplitude prediction, were determined through a hyperparameter search. We conjecture that a smaller $\rho$ is optimal for amplitude scale prediction since this leads to representations that are especially strong at discriminating amplitude differences among small groups of sensors. Perhaps this makes it easier to distinguish between neural responses from distinct parts of the brain such as the STG, which is associated with speech onset (Hamilton et al., 2018). In contrast, a larger $\rho$ for phase shift prediction could lead to representations that better discriminate neural synchrony information which is distributed across the brain rather than localised. As a result, a large proportion of the sensors in a MEG scanner should encode information about this feature.
@@ 4.3 Scaling Speech Decoding With Unlabelled Data
Figure 3: Scaling unlabelled data improves generalisation. We pre-train the model on increasing amounts of unlabelled data from Cam-CAN (Shafto et al., 2014; Taylor et al., 2017). The solid lines are the best linear fits to the data and the dashed lines show the amount of data used in prior surgical (Moses et al., 2021) and non-invasive (Défossez et al., 2023) work.
@@
Here, we analyse generalisation as we increase the volume of unlabelled data, analysing scaling performance on downstream tasks. As before, we pre-train with the combined pretext tasks. Figure 3 shows ROC AUC as we increase the amount of unlabelled data in pre-training up to approximately 160 hours. For both tasks, pre-training with any amount of data is sufficient to beat chance and there is a clear improvement in accuracy as the amount of unlabelled data increases. For speech detection on Armeni et al. (2022), scaling appears logarithmic in log-space; for all others, ROC AUC improves log-linearly within the data regime we study. In any case, adding unlabelled data has improved generalisation. Notably, we have scaled far beyond the data regime of prior surgical and non-surgical work and yet performance has continued to scale. Thus, our self-supervision approach may remain useful as the volume of open data in the field continues to rapidly increase.
@@
Our results also reveal several new and notable phenomena. Firstly, we scaled up the pre-training dataset by increasing the number of subjects. Since this led to consistent and almost monotonic improvements in downstream accuracy, our method is an exception to the common consensus that pooling subjects worsens generalisation. Secondly, as we pre-trained our model with a different dataset to those we fine-tuned on, our representation shows cross-dataset generalisation. This is particularly surprising as the Armeni et al. (2022), Gwilliams et al. (2023), and our pre-training dataset all use different scanners entirely. Performing well across these datasets indicates that, together, our architecture and pretext tasks successfully generate representations that are generalisable across heterogeneous scanners. Finally, we note that our pre-training dataset contained no language data whatsoever yet still improved downstream accuracy on language tasks. Remarkably, this shows that unlabelled brain data collected from any task (including those that are not linguistic) can be used to improve speech decoding performance.
@@
Since the results show improvements on both downstream tasks, this indicates that our pretext tasks are sufficiently generic to produce representations that work with multiple speech decoding tasks while still generalising well on each task individually. This is generally a challenging trade-off to manage. However, we notice that in both tasks, the base accuracy is higher and the improvement in ROC AUC is steeper for Armeni et al. (2022). This is likely to be because this dataset has more within-subject data. The weaker results for Gwilliams et al. (2023) may be a consequence of the larger number of subjects with shorter intra-subject recordings and greater subject variation. These observations support the findings of other recent work such as Csaky et al. (2022).
@@ 4.4 Scaling Unlabelled Data Improves Generalisation To Novel Subjects
In neuroimaging, brain data is generally highly variable across participants, leading to difficulty transferring models to novel subjects (Csaky et al., 2022). Whilst we have shown generalisation across subjects, here, we investigate whether we can generalise to novel subjects—an even more difficult challenge. This is critical in order to widely deploy speech BCIs for new patients. In this experiment, we fine-tune only on Gwilliams et al. (2023) and hold out three subjects with which we evaluate novel subject generalisation.
Figure 4: Scaling unlabelled data improves novel subject generalisation. We fine-tune on Gwilliams et al. (2023). When in-distribution, we evaluate on held-out sessions from subjects in the training set; when out-of-distribution, we evaluate on three held-out subjects. The solid lines are the best linear fits while the dashed lines show the amount of data used in prior surgical (Moses et al., 2021) and non-invasive (Défossez et al., 2023) work.
Figure 4 shows that scaling up the amount of unlabelled data used in pre-training not only improves accuracy on subjects previously seen, but also demonstrates a positive log-linear trend in performance for novel subjects. This indicates that scaling our method is an encouraging direction for resolving the challenges of subject variance faced by prior work. Moreover, as far as we are aware, this is the first result to demonstrate novel subject generalisation in speech decoding from MEG.
@@ 4.5 Aggregating Unlabelled MEG Datasets
To scale up unlabelled data further than individual studies, we must be able to combine many existing datasets. As a preliminary investigation, we combine two of the largest public MEG datasets: MOUS (Schoffelen et al., 2019) and Cam-CAN (Shafto et al., 2014; Taylor et al., 2017). In this section, we investigate how pre-training with these combined datasets affects downstream performance using the same experimental setup as Figure 3.
Pre-training Armeni Gwilliams
dataset Hours ROC AUC $t$ $p$ ROC AUC $t$ $p$
Cam-CAN 159 $\mathbf{0.621}\pm 0.003$ $36$ $4\mathrm{e}{-4}$ $0.543\pm 0.003$ $13$ $3\mathrm{e}{-3}$
MOUS 160 $0.605\pm 0.000$ $261$ $7\mathrm{e}{-6}$ $0.543\pm 0.004$ $9$ $5\mathrm{e}{-3}$
Cam-CAN + MOUS 319 $0.611\pm 0.003$ $40$ $3\mathrm{e}{-4}$ $\mathbf{0.546}\pm 0.002$ $20$ $1\mathrm{e}{-3}$
Table 2: Combining unlabelled datasets shows signs of outperforming single studies. We examine performance on the speech detection task. We see a small improvement when the datasets are combined on Gwilliams et al. (2023), but not Armeni et al. (2022).
The results in Table 2 show, for the first time, that combining datasets can improve performance on downstream speech decoding tasks. It leads to better performance on Gwilliams et al. (2023) compared to pre-training on either dataset alone. Interestingly, this was not the case for Armeni et al. (2022) where pre-training on Cam-CAN alone performed best. Combined pre-training did, however, outperform training only on MOUS. It is surprising that pre-training on Cam-CAN was better than pre-training on MOUS when evaluating on Armeni et al. (2022) given that MOUS and Armeni et al. (2022) both used speech tasks and were acquired on the same MEG scanner. Cam-CAN, by contrast, did not use a speech task and was acquired on a different MEG scanner. We hypothesise that the better results for Cam-CAN are due to it being a cleaner dataset. During our experiments, we found that data quality, even among unlabelled data, can have a significant affect as artefacts in recordings disrupt learning.
While the combination of the two datasets includes far more hours of data than any prior work on deep learning with MEG, further work needs to be done to aggregate more datasets. Here, we were limited by compute budget. Increasing the number of datasets could enable the network to eventually always improve over the best singular dataset. Just as increasing the number of subjects (rather than only within-subject data) improves novel subject generalisation, a larger number of datasets may be key to scaling results when datasets are aggregated in pre-training.

### v4

@@ 4 Experiments
We evaluate our self-supervised representations by measuring how they scale with unlabelled data and generalise across datasets, subjects, and tasks. We focus our evaluation on MEG data as the signal is rich, with better spatial resolution than EEG (Lopes da Silva, 2013) and faster sampling rates than fMRI (Hall et al., 2014). We pre-train all models to completion and then train a linear probe on labelled data for each task. In all tables and figures, we quote the receiver operating characteristic area under the curve (ROC AUC) where chance is always 0.5 regardless of the class distribution. We show the test ROC AUC at the best validation ROC AUC (early stopping) and quote uncertainty as the standard error of the mean over up to five seeds.
@@ 4.1 Experimental setup
Datasets. In total, we use almost five times the volume of data in prior MEG work, totalling approximately 400 hours with nearly 900 subjects across pre-training and downstream training. Unless specified otherwise, we pre-train with Cam-CAN (Shafto et al., 2014; Taylor et al., 2017) as an unlabelled representation learning dataset. This is a study containing 641 subjects with resting and sensorimotor tasks, totalling approximately 160 hours of MEG recordings. When aggregating datasets, we also pre-train with MOUS (Schoffelen et al., 2019), which contains 204 subjects and another 160 hours from visual and auditory tasks. Downstream, we use labelled heard speech MEG datasets where participants listen to short stories or audiobooks. We mainly focus on Armeni et al. (2022) which contains 3 subjects who listen to 10 hours of recordings each (30 hours total). We also analyse Gwilliams et al. (2023) which has 27 subjects, each recorded for 2 hours (54 hours total). The latter dataset is particularly difficult to decode from as there is very little within-subject data and it did not enforce the use of head casts to immobilise participants. Nevertheless, given it has many more subjects, we use this dataset to study subject generalisation.
Preprocessing. Each recording is in $\mathbb{R}^{S\times T}$ where $S$ is the number of sensors and $T$ is the number of time points sampled by the scanner. To eliminate high-frequency artifacts, we apply a low-pass filter at 125Hz as well as a high-pass filter at 0.1Hz to remove slow-drift artifacts. Since the datasets were recorded in Europe, where the electric grid frequency is 50Hz, we apply a notch filter at 50Hz and its harmonics to account for line noise. Treating the low-pass filter threshold as the Nyquist frequency, we downsample the signal to twice that at 250Hz, avoiding aliasing within our band of interest. Finally, we detect sensor channels with significant noise and artifacts using a variance threshold and replace them by interpolating the spatially nearest sensors.
Downstream tasks. We evaluate our methods primarily on speech detection. This task determines whether speech occurs in the auditory stimulus using the neural response. This is a fundamental task in understanding speech perception and is one of the few tasks that so far show statistically significant results in highly noisy MEG signals. It also has direct applications to BCIs as it can be used to segment words or sentences for decoding and activate a speech BCI when a patient wishes to communicate. Secondarily, we also study voicing classification to demonstrate the versatility of our representations for general speech decoding tasks. Given data aligned at the onset of a phoneme, the task is to recognise whether the phoneme is voiced or voiceless, where voicing is a phonetic feature that categorises whether a speech sound is associated with vocal cord vibration. This task is also directly relevant to a speech BCI as it involves classifying phonemes which can be used to decode words.
@@ 4.2 Learning Generalisable Representations Using Pretext Tasks
Table 2 shows that our approach achieves two key feats: outperforming comparable state-of-the-art self-supervised methods by 15-27% (part C), and matching the performance of prior self-supervised methods with surgical data (11) while using only non-invasive data. Since non-invasive data has a much lower signal-to-noise ratio than surgical data, this is quite unprecedented. In the rest of this section, we analyse this table in detail.
@@
In part B, we show the results of pre-training models with each pretext task independently, together, and without any pre-training at all. Using pretext tasks (3, 4, 5) outperforms no pre-training (2). Interestingly, the combination of all pretext tasks (5) leads to better generalisation than any task on its own (the improvement over (3) is statistically significant). We conjecture that this is because our pretext tasks capture complementary properties in time- and frequency-space, ensuring that the representation includes more salient features than any individual task could encode. Finally, we apply Gaussian filtering to the predictions (7), smoothing out anomalies in the predicted speech envelope.
Part / ID Model ROC AUC
A 0 Random select .500.500.500.500
1 Linear .$539_{\pm.002}.$
B 2 Ours No pre-train. .$519_{\pm.002}.$
3 + linear Amp($\rho=0.2$) .$624_{\pm.001}.$
4 Phase($\rho=0.5$) .$615_{\pm.001}.$
5 Band .$588_{\pm.001}.$
6 All tasks .$630_{\pm.000}.$
7  + smoothing .$\mathbf{.700}_{\pm.002}$
C* 8 BIOT1 + linear .$615_{\pm.002}.$
9 BrainBERT2 + linear .$556_{\pm.007}.$
10 Ours (best) + linear .$\mathbf{.705}_{\pm.003}$
11 BrainBERT2 + lin. (surgical) .$71_{\pm.06}.$
1Yang et al. (2023) 2Wang et al. (2023)
@@
Table 2: Our approach outperforms all baselines and matches surgical decoding. In the linear-only case, we train a supervised linear classifier on the input MEG signals. For ours and BrainBERT, we train a linear layer on top of a backbone pre-trained on CamCAN, with the rest of the model frozen. For BIOT, we use their pre-trained weights. In the no pre-training baseline, the backbone uses randomly initialised and frozen weights. In the surgical context, we quote the relevant result from (Wang et al., 2023, Table 2). With all pretext tasks, their losses are weighted equally.
Next, we turn to the baselines, starting with Table 2 part A. Our method outperforms random selection (0) and training a linear layer with the MEG signal directly (1). The latter even has substantially more trainable parameters because the input dimension is larger without an encoder. In part C, we compare our approach to two state-of-the-art self-supervised methods. In each experiment, we apply Gaussian filtering as in (7). Although better than random, BrainBERT (9) does not generalise as well as our method (10). BrainBERT employs a generic masked spectrogram in-filling pre-training objective. While it is intended for speech decoding tasks, the pre-training objective is not designed to specifically capture features salient to neural speech processing. Furthermore, their method is less data-efficient because in-filling is a harder generative task compared to classification and while their methodology was developed for relatively high-fidelity intracranial recordings, the inherently lower signal-to-noise ratio of MEG presents an even greater challenge.
Our last baseline is BIOT, which we were unable to train successfully beyond chance level using MEG data. Instead, in (8), we leverage publicly released weights pre-trained with over fifty thousand hours of EEG. However, BIOT still falls short of our pre-training method for reasons which we believe are similar to BrainBERT—its objective does not leverage neuroscientific understanding of speech processing.
Finally, and quite remarkably, our best result matches the AUROC quoted in Wang et al. (2023, Table 2) who use intracranial data from heard speech (11). We achieved this score with non-invasive data which is typically substantially more difficult to decode due to the low signal-to-noise ratio.
00footnotetext: Due to the large computational cost of processing embeddings for MEG data in BrainBERT, we restrict pre-training of experiments in part C to approximately 30 hours and use only subject 001 of the downstream dataset for training and evaluation.
@@ 4.3 Scaling Speech Decoding With Unlabelled Data
Figure 3: Scaling unlabelled data improves generalisation. We pre-train the model on increasing amounts of unlabelled data from Cam-CAN (Shafto et al., 2014; Taylor et al., 2017). The solid lines are linear fits and the dashed lines show the volume of data used in prior surgical (Moses et al., 2021) and non-invasive (Défossez et al., 2023) work. Unlike Table 2 (7) we do not apply Gaussian filtering to the predictions for simplicity. The best improvements are statistically significant.
@@
Figure 3 shows that performance scales predictably with unlabelled data volume, following distinct patterns for different tasks and datasets. For speech detection on Armeni et al. (2022), we observe logarithmic scaling in log-space (log-log scaling), suggesting diminishing but continued returns with increased data. For other tasks, ROC AUC improves log-linearly, indicating robust scaling potential. Importantly, even our smallest pre-training dataset beats chance performance, while our largest (160 hours) continues to show gains without plateauing. Notably, we have scaled far beyond the data regime of prior surgical and non-surgical work and yet performance has continued to scale. Thus, our self-supervision approach may remain useful as the volume of open data in the field continues to rapidly increase.
@@
Our results also reveal several new and notable phenomena. We scaled up the pre-training dataset by increasing the number of subjects. Since this led to consistent and almost monotonic improvements in downstream accuracy, our method is an exception to the consensus that pooling subjects worsens generalisation. As we pre-trained our model with a different dataset to those we fine-tuned on, our representation shows cross-dataset generalisation. This is surprising as the Armeni et al. (2022), Gwilliams et al. (2023), and our pre-training dataset all use different scanners entirely. Performing well across these datasets indicates that, together, our architecture and pretext tasks successfully generate representations that are generalisable across heterogeneous scanners. Finally, we note that our pre-training dataset contained no language data whatsoever yet still improved downstream accuracy on language tasks. Remarkably, this shows that unlabelled brain data collected from any task (including those that are not linguistic) can be used to improve speech decoding performance.
@@
Since the results show improvements on both downstream tasks, this indicates that our pretext tasks are sufficiently generic to produce representations that work with multiple speech decoding tasks while still generalising well on each task individually. This is generally a challenging trade-off to manage. However, we notice that in both tasks, the base accuracy is higher and the improvement in ROC AUC is steeper for Armeni et al. (2022). This is likely to be because this dataset has more within-subject data. The weaker results for Gwilliams et al. (2023) may be a consequence of shorter intra-subject recordings, greater subject variation, and the lack of head casts in data collection. These observations support the findings of work such as Csaky et al. (2022).
@@ 4.4 Scaling Unlabelled Data Improves Generalisation To Novel Subjects
In neuroimaging, brain data is generally highly variable across participants, leading to difficulty transferring models to novel subjects (Csaky et al., 2022). Whilst we have shown generalisation across subjects, here, we investigate whether we can generalise to novel subjects—an even more difficult challenge. This is critical in order to widely deploy speech BCIs for new patients. In this experiment, we use Gwilliams et al. (2023) as our downstream dataset because of the large number of participants, holding out three subjects with which we evaluate novel subject generalisation.
Figure 4: Scaling unlabelled data improves novel subject generalisation. We train a linear probe on Gwilliams et al. (2023). When in-distribution, we evaluate on held-out sessions; when out-of-distribution, we evaluate on held-out subjects. The lines represent the same as in Figure 3. The best improvements are statistically significant.
Figure 4 shows that scaling up the amount of unlabelled data used in pre-training not only improves accuracy on subjects previously seen, but also demonstrates a positive log-linear trend in performance for novel subjects. This indicates that scaling our method is an encouraging direction for resolving the challenges of subject variance faced by prior work. As far as we are aware, this is the first result to demonstrate novel subject generalisation in speech decoding from MEG.
@@ 4.5 Aggregating Unlabelled MEG Datasets
Given the promising scaling results with single datasets, a natural question arises: can we achieve even better performance by combining multiple MEG datasets? This is particularly challenging since datasets often use different scanning hardware and experimental protocols. Thus, it has so far not been shown in MEG.
As a preliminary investigation, we combine two of the largest public MEG datasets: MOUS (Schoffelen et al., 2019) and Cam-CAN (Shafto et al., 2014; Taylor et al., 2017). In this section, we investigate how pre-training with these combined datasets affects downstream performance using the same experimental setup as Figure 3.
Pre-training Data Hours ROC AUC
CamCAN1,2 $159$$630_{\pm.0001}.$
MOUS3 $160$$614_{\pm.0004}.$
CamCAN1,2 + MOUS3 $319$$\mathbf{.638}_{\pm.0002}$
1Shafto et al. (2014) 2Taylor et al. (2017)
3Schoffelen et al. (2019)
Table 3: Aggregating unlabelled datasets outperforms single studies. For the first time with MEG data, we show that unlabelled pre-training data from multiple studies with different hardware profiles can be aggregated while continuing to gain the benefits of scaling. Combining data leads to a statistically significant $(p<0.05)$ improvement.
The results in Table 3 show, for the first time, that combining datasets can improve performance on downstream speech decoding tasks. It leads to better performance compared to pre-training on either dataset alone. It is surprising that pre-training on Cam-CAN was better than pre-training on MOUS given that MOUS and the downstream dataset both used speech tasks and were acquired on the same MEG scanner. Cam-CAN, by contrast, did not use a speech task and was acquired on a different MEG scanner. We hypothesise that the better results for Cam-CAN are due to it being cleaner. During our experiments, we found that data quality, even among unlabelled data, can have a significant effect as artefacts in recordings disrupt learning.
While the combination of the two datasets includes far more hours of data than any prior work on deep learning with MEG, further work needs to be done to aggregate more datasets. Here, we were limited by compute budget and data availability. Increasing the number of datasets (e.g., by including EEG too) could further improve results. Just as increasing the number of subjects (rather than only within-subject data) improves novel subject generalisation, a larger number of datasets may be key to scaling results when datasets are aggregated in pre-training.